# Accelerating Goal-Conditioned Reinforcement Learning Algorithms and Research

### v2

@@ Abstract
Self-supervision has the potential to transform reinforcement learning (RL), paralleling the breakthroughs it has enabled in other areas of machine learning. While self-supervised learning in other domains aims to find patterns in a fixed dataset, self-supervised goal-conditioneds reinforcement learning (GCRL) agents discover new behaviors by learning from the goals achieved during unstructured interaction with the environment. However, these methods have failed to see similar success, both due to a lack of data from slow environment simulations as well as a lack of stable algorithms. We take a step toward addressing both of these issues by releasing a high-performance codebase and benchmark (JaxGCRL) for self-supervised GCRL, enabling researchers to train agents for millions of environment steps in minutes on a single GPU. By utilizing GPU-accelerated replay buffers, environments, and a stable contrastive RL algorithm, we reduce training time by up to $\times2$. Additionally, we assess key design choices in contrastive RL, identifying those that most effectively stabilize and enhance training performance. With this approach, we provide a foundation for future research in self-supervised GCRL, enabling researchers to quickly iterate on new ideas and evaluate them in diverse and challenging environments. Code: https://github.com/MichalBortkiewicz/JaxGCRL
@@ 1 Introduction
Self-supervised learning has significantly influenced machine learning over the last decade, transforming how research is done in domains such as natural language processing and computer vision (Chen et al., 2020b, ; Dosovitskiy et al.,, 2020; Vaswani et al.,, 2017). In the context of reinforcement learning (RL), most self-supervised prior methods apply the same recipe that has been successful in other domains: learning representations or models from a large, fixed dataset (Hoffmann et al.,, 2022; Sardana et al.,, 2024; Muennighoff et al.,, 2023). However, the RL setting also enables a fundamentally different type of self-supervised learning: rather than learning from a fixed dataset (as done in NLP and computer vision), a self-supervised reinforcement learner can collect its own dataset. Thus, rather than learning a representation of a dataset, the self-supervised reinforcement learner acquires a representation of an environment or of behaviors and optimal policies therein. Self-supervised reinforcement learners may address many of the challenges that stymie today’s foundations models: reasoning about the consequences of actions (Rajani et al.,, 2019; Bhargava and Ng,, 2022; Kwon et al.,, 2023) (i.e., counterfactuals (Valmeekam et al.,, 2022; Jin et al.,, 2023)) and long horizon planning (Valmeekam et al.,, 2022; Du et al.,, 2023; Guan et al.,, 2023).
@@
The main goal of this paper is to introduce JaxGCRL: extremely fast GPU-accelerated codebase and benchmark for effective self-supervised GCRL research. For instance, an experiment with $10$ million environment steps lasts only around 10 minutes on a single GPU, which is up to 22$\times$ faster than in the original contrastive RL codebase (Eysenbach et al.,, 2022) (Fig. 1). This speed allows researchers to “interactively” debug their algorithms, changing pieces and getting results in near real-time for multiple random seeds on a single GPU without the hustle of distributed training. Consequently, JaxGCRL eliminates the barriers to entry to state-of-the-art GCRL research, making it more accessible to under-resourced institutions.
@@ 2.2 Accelerating Deep Reinforcement Learning
Deep RL has only recently become practical for many tasks, in part due to improvements in hardware support for these algorithms. Distributed training has enabled RL algorithms to scale across hundreds of GPUs (Mnih et al.,, 2016; Espeholt et al.,, 2018; 2020; Hoffman et al.,, 2022). To resolve the bottleneck of environment interaction with CPU-bound environments, various GPU-accelerated environments have been proposed (Matthews et al.,, 2024; Freeman et al.,, 2021; Bonnet et al.,, 2024; Rutherford et al.,, 2024; Liang et al.,, 2018; Dalton et al.,, 2020; Makoviychuk et al.,, 2021). In addition to reducing CPU-GPU data-transfer overhead, these environments can easily be vectorized to collect data from hundreds of rollouts in parallel (Bradbury et al.,, 2018; Heek et al.,, 2023; Hennigan et al.,, 2020). We build on these advances to scale self-supervised RL to data-rich settings.
@@ 2.4 RL Benchmarks
The RL community has recently started to pay greater attention to how RL research is conducted, reported, and evaluated (Henderson et al.,, 2018; Jordan et al.,, 2020; Agarwal et al.,, 2022). A key issue is a lack of reliable and efficient benchmarks: it becomes hard to rigorously compare novel methods when the number of trials needed to see statistically significant results across diverse settings ranges in the thousands of training hours (Jordan et al.,, 2024). Some benchmarks that have seen adoption include OpenAI gym/Gymnasium (Brockman et al.,, 2022; Towers et al.,, 2024), DeepMind Control Suite (Tassa et al.,, 2018), and D4RL (Fu et al.,, 2020). More recently, hardware-accelerated versions of some of these benchmarks have been proposed (Gu et al.,, 2021; Matthews et al.,, 2024; Koyamada et al.,, 2023; Makoviychuk et al.,, 2021; Nikulin et al.,, 2023). However, the RL community still lacks advanced benchmarks for goal-conditioned methods. We address this gap with JaxGCRL, significantly lowering the GCRL evaluation cost and thereby enabling impactful RL research.
@@ 3.2 Policy Learning
To form a batch of data, we sample uniformly $(s,a)$ pairs and then sample goals $g$ from states that occur after $s$ in a trajectory.
@@ 5.1 Experimental Setup
Our experiments use JaxGCRL suite of simulated environments described in Section 4.2. We evaluate algorithms in an online setting, with a UTD ratio $1$:$16$, using a batch size of $256$ and a discount factor of $0.99$. For every environment, we sample evaluation goals from the same distribution as training ones and use a replay buffer of size $10$M. We use $1024$ parallel environments to gather the data and run experiments for $50$M environment steps. We compare CRL with Soft Actor-Critic (SAC) (Haarnoja et al.,, 2018) and SAC with Hindsight Experience Replay (HER) (Andrychowicz et al.,, 2017). For a fair comparison, we use a sparse reward for both SAC and SAC+HER baselines, with $r=1$ when the agent is in goal proximity and $r=0$ otherwise. For SAC HER, we use final strategy relabelling, i.e. relabeling goals with states achieved at the end of the trajectory. In the majority of experiments, we use CRL with L2 energy function, symmetric InfoNCE objective, and a tuneable entropy coefficient for all methods, see Appendix B. We use two performance metrics: success rate and time near goal. Success rate measures whether the agent reached the goal at least once during the episode, while time near goal indicates how long the agent stayed close to it.
@@ 5.2 JaxGCRL benchmark results
We establish baseline results for JaxGCRL with CRL, SAC and SAC+HER algorithms in Fig. 3. Clearly, CRL achieves the highest performance across tested methods, resulting in non-trivial policies even in the hardest tasks in terms of high-dimensional state and action spaces (Humanoid) and exploration (Ant Pusher and Ant Push). However, the performance in these challenging environments is low, indicating room for improvement for future contrastive RL methods. Unsurprisingly, SAC+HER results in higher performance than SAC, with one exception for the Pusher environment where HER is generating a trivial relabeled experience; please refer to Section A.5 for a detailed explanation. Additional experiments on JaxGCRL environments with design choices discussed in the following sections can be found in Section A.4.
Figure 3: Baseline results in JaxGCRL benchmark. Success rates of CRL, SAC, and SAC+HER for $50$M environment steps for each JaxGCRL environments. CRL outperforms other baselines in most of the environments. The training speed is a function of the environment complexity, method complexity, and physics backend; see Section A.4. Results are reported as the interquartile mean (IQM) along with its standard error, based on 10 seeds.
@@ 5.3 Contrastive objectives and energy functions comparison
Contrastive objective and energy function are the two main components of contrastive methods, serving as the primary drivers of their final performance. We evaluate CRL with 10 different contrastive objectives, as defined in Section 4.3 across five environments: Ant Soccer, Ant, Ant Big Maze, Ant U-Maze and Pusher, and report aggregated performance. For the energy function, we use L2, as it consistently resulted in the highest performance for the CRL method, especially regarding time near goal, see Section A.2. Additionally, we apply logsumexp regularization with a coefficient of 0.1 to each objective. This auxiliary objective is essential, as without it, the performance of InfoNCE deteriorates significantly (Eysenbach et al.,, 2022). The analysis presented in Fig. 4 reveals that the originally proposed NCE-binary objective, along with forward-backward, IPO, and SPPO, are the least effective objectives among those evaluated. However, for other InfoNCE-derived objectives, it is difficult to determine the best one, as their performance is similar. Interestingly, CRL seems fairly robust to the choice of contrastive objective.
Figure 4: InfoNCE-based loss functions perform best. The critic loss functions that result in the highest success rates are those based on InfoNCE and DPO, though DPO policies remain at the goal for a shorter. IQMs averaged over 10 seeds and plotted with one standard error.
@@ 5.5 Scaling the data
We evaluate the benefits CRL gains from training in a data-rich setting. In particular, we report performance for large architectures (studied in Section 5.4) with different combinations of energy functions and contrastive objectives for $300$M environment steps in Fig. 8. We observe that the L2 energy function with InfoNCE objective configuration outperforms all others by a substantial margin, leading to a higher success rate and time near the goal across three locomotion tasks. Interestingly, the dot product energy function performs best in the object manipulation task (Ant Soccer). This indicates that only a subset of a wide array of design choices performs well when scaling CRL. Additionally, there is still room for improvement in scaling CRL with data, as the success rate in Ant Soccer and Ant Big Maze remains around $40\%$.
@@ 5.6 Gradient updates to data ratio
To find a good balance between CRL performance and training speed, we evaluate different ratios of update to data. In particular, we examine ratios ($1$:$1$, $1$:$8$, $1$:$16$, $1$:$24$, $1$:$32$, $1$:$49$, $1$:$48$) in four environments: Ant, Ant Soccer, Ant U-Maze and Pusher Hard. Interestingly, we only observe an increase in performance for Pusher Hard with a higher number of updates, while in other environments, it leads to decreased performance. With UTD ratio of $1:16$, our code is $\times2$ faster than prior implementations and with a lower frequency of gradient updates can be even faster.
@@
Experiments with 10M steps can be completed in minutes, while those with billions of environment steps can be done in a few hours using JaxGCRL on a single GPU. CRL is the only method that can learn effectively in all proposed environments without needing a high UTD ratio in most cases. It benefits greatly from using large architectures, especially when Layer Normalization is applied. Different combinations of energy and contrastive functions lead to different outcomes: some primarily improve success rates, while others extend the time spent near the goal.
@@ 6 Conclusion
All experiments can be replicated using the provided publicly available JaxGCRL code at https://anonymous.4open.science/r/JaxGCRL-2316/README.md. This repository includes comprehensive instructions for setting up the environment, running the experiments, and evaluating the results, making it straightforward to reproduce the findings.
@@ B.1 Environment details
B.2 Benchmark parameters
@@ Appendix F Pseudocode
2:Initialize $\phi$, $\psi$, $\pi$ and an empty replay buffer $\mathcal{D}$
3:repeat
4:     In parallel for each environment instance:
5:         Observe state $s$ and sample an action $a\sim\pi(s,g)$
6:         Execute $a$ in the environment
7:         Observe next state $s^{\prime}$ and done signal $d$ to indicate whether $s^{\prime}$ is terminal
8:         Append $(s,a,s^{\prime})$ to current trajectory for this environment
9:         if $s^{\prime}$ is terminal then
10:              Reset this environment state (this includes sampling new goal)
11:              Store current trajectory for this environment in $\mathcal{D}$
12:              Start new trajectory for this environment
13:     if it is time to update then
14:         for $j$ in range(number of updates) do
15:              From buffer $\mathcal{D}$ randomly sample (with discount) a batch $\mathcal{B}$ of state-action pairs and goals from their future
16:              Update critic by one step of gradient descent using given Critic loss:
$\nabla_{\phi,\psi}\,\mathcal{L}_{\text{Critic}}(\mathcal{B};\phi,\psi)+\beta\mathcal{L}_{\text{logsumexp}}(\mathcal{B},\phi,\psi)$
17:              Update policy by one step of gradient descent using:
$\nabla_{\pi}\,\mathcal{L}_{\text{Actor}}(\mathcal{B};\phi,\psi,\pi)$
18:until convergence

### v3

@@ Abstract
Self-supervision has the potential to transform reinforcement learning (RL), paralleling the breakthroughs it has enabled in other areas of machine learning. While self-supervised learning in other domains aims to find patterns in a fixed dataset, self-supervised goal-conditioneds reinforcement learning (GCRL) agents discover new behaviors by learning from the goals achieved during unstructured interaction with the environment. However, these methods have failed to see similar success, both due to a lack of data from slow environment simulations as well as a lack of stable algorithms. We take a step toward addressing both of these issues by releasing a high-performance codebase and benchmark (JaxGCRL) for self-supervised GCRL, enabling researchers to train agents for millions of environment steps in minutes on a single GPU. By utilizing GPU-accelerated replay buffers, environments, and a stable contrastive RL algorithm, we reduce training time by up to $\times2$. Additionally, we assess key design choices in contrastive RL, identifying those that most effectively stabilize and enhance training performance. With this approach, we provide a foundation for future research in self-supervised GCRL, enabling researchers to quickly iterate on new ideas and evaluate them in diverse and challenging environments. Code: https://github.com/MichalBortkiewicz/JaxGCRL.
@@ 1 Introduction
Self-supervised learning has significantly influenced machine learning over the last decade, transforming how research is done in domains such as natural language processing and computer vision (Chen et al., 2020b, ; Dosovitskiy et al.,, 2020; Vaswani et al.,, 2017). In the context of reinforcement learning (RL), most self-supervised prior methods apply the same recipe that has been successful in other domains: learning representations or models from a large, fixed dataset (Hoffmann et al.,, 2022; Sardana et al.,, 2024; Muennighoff et al.,, 2023). However, the RL setting also enables a fundamentally different type of self-supervised learning: rather than learning from a fixed dataset (as done in NLP and computer vision), a self-supervised reinforcement learner can collect its own dataset. Thus, rather than learning a representation of a dataset, the self-supervised reinforcement learner acquires a representation of an environment or of behaviors and optimal policies therein. Self-supervised reinforcement learners may address many of the challenges that stymie today’s foundation’s models: reasoning about the consequences of actions (Rajani et al.,, 2019; Kwon et al.,, 2023) (i.e., counterfactuals (Bhargava and Ng,, 2022; Jin et al.,, 2023)) and long horizon planning (Bhargava and Ng,, 2022; Du et al.,, 2023; Guan et al.,, 2023).
@@
The main goal of this paper is to introduce JaxGCRL: an extremely fast GPU-accelerated codebase and benchmark for effective self-supervised GCRL research. For instance, an experiment with $10$ million environment steps lasts only around 10 minutes on a single GPU, which is up to 22$\times$ faster than in the original contrastive RL codebase (Eysenbach et al.,, 2022) (Fig. 1). This speed allows researchers to “interactively” debug their algorithms, changing pieces and getting results in near real-time for multiple random seeds on a single GPU without the hustle of distributed training. Consequently, JaxGCRL eliminates the barriers to entry to state-of-the-art GCRL research, making it more accessible to under-resourced institutions.
@@ 2.2 Accelerating Deep Reinforcement Learning
Deep RL has only recently become practical for many tasks, in part due to improvements in hardware support for these algorithms. Distributed training has enabled RL algorithms to scale across hundreds of GPUs (Mnih et al.,, 2016; Espeholt et al.,, 2018; 2020; Hoffman et al.,, 2022). To resolve the bottleneck of environment interaction with CPU-bound environments, various GPU-accelerated environments have been proposed (Matthews et al.,, 2024; Freeman et al.,, 2021; Bonnet et al.,, 2024; Rutherford et al.,, 2024; Liang et al.,, 2018; Dalton et al.,, 2020; Makoviychuk et al.,, 2021; Lange,, 2022; Lu et al.,, 2022). Most of these works rely on JAX (Bradbury et al.,, 2018; Heek et al.,, 2023; Hennigan et al.,, 2020), which enables JIT compilation, operator fusion and other components necessary for efficient vectorized code execution. These features significantly accelerate data collection by supporting rollouts in hundreds of parallelized environments. We build on these advances to scale self-supervised RL to data-rich settings.
@@ 2.4 RL Benchmarks
The RL community has recently started to pay greater attention to how RL research is conducted, reported, and evaluated (Henderson et al.,, 2018; Jordan et al.,, 2020; Agarwal et al.,, 2022). A key issue is a lack of reliable and efficient benchmarks: it becomes hard to rigorously compare novel methods when the number of trials needed to see statistically significant results across diverse settings ranges in the thousands of training hours (Jordan et al.,, 2024). Some benchmarks that have seen adoption include OpenAI gym/Gymnasium (Brockman et al.,, 2022; Towers et al.,, 2024), DeepMind Control Suite (Tassa et al.,, 2018), and D4RL (Fu et al.,, 2021). More recently, hardware-accelerated versions of some of these benchmarks have been proposed (Gu et al.,, 2021; Matthews et al.,, 2024; Koyamada et al.,, 2023; Makoviychuk et al.,, 2021; Nikulin et al.,, 2023). However, the RL community still lacks advanced benchmarks for goal-conditioned methods. We address this gap with JaxGCRL, significantly lowering the GCRL evaluation cost and thereby enabling impactful RL research.
@@ 3.2 Policy Learning
Each batch is formed by sampling $(s,a)$ pairs uniformly and then sampling goals $g$ from the states that occur after $s$ in a trajectory.
@@ 5.1 Experimental Setup
Our experiments use JaxGCRL suite of simulated environments described in Section 4.2. We evaluate algorithms in an online setting for $50$M environment steps. We compare CRL with Soft Actor-Critic (SAC) (Haarnoja et al.,, 2018), SAC with Hindsight Experience Replay (HER) (Andrychowicz et al.,, 2017), TD3 (Fujimoto et al.,, 2018), TD3+HER, and PPO (Schulman et al.,, 2017). For algorithms with HER, we use final strategy relabelling, i.e. relabeling goals with states achieved at the end of the trajectory. In the majority of experiments, we use CRL with L2 energy function, symmetric InfoNCE objective, and a tuneable entropy coefficient for all methods. See Appendix B for details. We use two performance metrics: success rate and time near goal. Success rate measures whether the agent reached the goal at least once during the episode, while time near goal indicates how long the agent stayed close to it. We use a sparse reward for all baselines, with $r=1$ when the agent is in goal proximity and $r=0$ otherwise. We define goal-reaching as achieving proximity below the goal distance threshold defined in Table 1. The implementations of PPO and SAC are sourced from the Brax repository (Freeman et al.,, 2021), while TD3, HER, and CRL are partially based on Brax.
@@ 5.2 JaxGCRL benchmark results
We establish baseline results for JaxGCRL with all algorithms in Fig. 3. Clearly, CRL achieves the highest performance across tested methods, resulting in non-trivial policies even in the hardest tasks in terms of high-dimensional state and action spaces (Humanoid) and exploration (Pusher and Ant Push). However, the performance in these challenging environments is low, indicating room for improvement for future contrastive RL methods. As expected, HER can improve performance for both TD3 and SAC. In contrast, PPO performs poorly across all tasks, likely due to the challenges posed by the sparse reward setting. Additional experiments on JaxGCRL environments with design choices discussed in the following sections can be found in Section A.4.
Figure 3: Baseline results in JaxGCRL benchmark. Success rates of all the baseline algorithms for $50$M environment steps for every JaxGCRL environment. CRL outperforms other baselines in most of the environments. The training speed is a function of the environment complexity, method complexity, and physics backend; see Section A.4. Specifically, due to differences in how each method works, the speed varies greatly in the same environments; this can be best seen with the PPO method being significantly faster than others due to it not using a replay buffer, which frees up GPU memory for more parallel environment simulations. Results are reported as the interquartile mean (IQM) along with its standard error, based on 10 seeds.
@@ 5.3 Contrastive objectives and energy functions comparison
Contrastive objective and energy function are the two main components of contrastive methods, serving as the primary drivers of their final performance. We evaluate CRL with 10 different contrastive objectives, as defined in Section 4.3 across five environments: Ant Soccer, Ant, Ant Big Maze, Ant U-Maze, Ant Push, and Pusher, and report aggregated performance. For the energy function, we use L2, as it consistently resulted in the highest performance for the CRL method, especially regarding time near goal, see Section A.2. Additionally, we apply logsumexp regularization with a coefficient of 0.1 to each objective. This auxiliary objective is essential, as without it, the performance of InfoNCE deteriorates significantly (Eysenbach et al.,, 2022). The analysis presented in Fig. 4 reveals that the originally proposed NCE-binary objective, along with forward-backward, IPO, and SPPO, are the least effective objectives among those evaluated. However, for other InfoNCE-derived objectives, it is difficult to determine the best one, as their performance is similar. Interestingly, CRL seems fairly robust to the choice of contrastive objective.
Figure 4: InfoNCE-based loss functions perform best. The critic loss functions that achieve the highest success rates are based on InfoNCE and DPO. However, DPO policies tend to stay at the goal for a shorter duration. IQMs averaged over 10 seeds and plotted with one standard error.
@@ 5.5 Scaling the data
We evaluate the benefits CRL gains from training in a data-rich setting. In particular, we report performance for large architectures (studied in Section 5.4) with different combinations of energy functions and contrastive objectives for $300$M environment steps in Fig. 8. We observe that the L2 energy function with InfoNCE objective configuration outperforms all others by a substantial margin, leading to a higher success rate and time near the goal across three locomotion tasks. Interestingly, the dot product energy function performs best in the object manipulation task (Ant Soccer). This indicates that only a subset of a wide array of design choices performs well when scaling CRL. Additionally, there is still room for improvement in scaling CRL with data, as the success rate in Ant Soccer and Ant Big Maze remains around $40\%$. For additional experiments, refer to Section A.6
@@ 5.6 Gradient updates to data ratio
JaxGCRL enables efficient execution of extensive experiments. Leveraging this capability, we explore the effect of the model’s update frequency in CRL by evaluating a range of UTD ratios. In particular, we examine ratios ($1$:$1$, $1$:$8$, $1$:$16$, $1$:$24$, $1$:$32$, $1$:$49$, $1$:$48$) in five environments: Ant, Ant Soccer, Ant U-Maze, Pusher Hard, and Ant Push. Interestingly, we only observe a significant increase in performance for Pusher Hard with a higher number of updates, while in other environments, it leads to decreased or similar performance. With a UTD ratio of $1$:$16$, our code is $\times2$ faster than prior implementations, and with a lower frequency of gradient updates, it can be even faster.
@@
  • Experiments with 10M steps can be completed in minutes, while those with billions of environment steps can be done in a few hours using JaxGCRL on a single GPU.
  • CRL is the only method that can learn effectively in all proposed environments without needing a high UTD ratio in most cases. It benefits greatly from using large architectures, especially when Layer Normalization is applied.
  • Different combinations of energy and contrastive functions lead to different outcomes: some primarily improve success rates, while others extend the time spent near the goal.
@@ 6 Conclusion
All experiments can be replicated using the provided publicly available JaxGCRL code at https://github.com/MichalBortkiewicz/JaxGCRL. This repository includes comprehensive instructions for setting up the environment, running the experiments, and evaluating the results, making it straightforward to reproduce the findings.
Acknowledgments.
This research was substantially supported by the National Science Centre, Poland (grant no. 2023/51/D/ST6/01609), and the Warsaw University of Technology through the Excellence Initiative: Research University (IDUB) program. We gratefully acknowledge the Polish high-performance computing infrastructure, PCSS PLCloud, for providing computational resources and support under grant no. pl0334-01, and PLGrid (HPC Center: ACK Cyfronet AGH), for providing resources and support under grant no. PLG/2024/017040. We would also like to recognize the founding of the DoD NDSEG fellowship for Vivek Myers. This work was partially conducted using Princeton Research Computing resources at Princeton University, a consortium led by the Princeton Institute for Computational Science and Engineering (PICSciE) and Research Computing.
@@ A.5 Hindsight experience replay details
A.6 Scaled-up CRL architecture in data-rich setting
In Figures 13 and 14, we report success rates and time near goal for scaled-up CRL agent with architecture consisting of $4$ layers with $1024$ neurons per layer and layer norm. We find that these architectures can increase the fraction of trials where the agent reaches the goal at least once, but they do not enable the agent to stabilise around the goal (e.g., on $7$ tasks, the best agent spends less than $50\%$ of an episode at the goal).
When visualizing the rollouts, we observe that the Humanoid agent falls immediately after reaching the goal state, and the Ant Soccer agent struggles to recover when it pushes the ball too far away. The Humanoid merely "flings" itself toward the goal, while the optimal policy would involve running to the goal and remaining there. This inability to stabilize around the goal suggests that the agent is not effectively optimizing the actor’s objective, pointing to a potential area for further research.
Figure 13: CRL with big architecture success rates in data-rich setting.
Figure 14: CRL with big architecture time near goal in data-rich setting.
@@ B.1 Environment details
B.2 Benchmark details
Our experiments use JaxGCRL suite of simulated environments described in Section 4.2. We evaluate algorithms in an online setting, with a UTD ratio $1$:$16$ for CRL, TD3, TD3+HER, SAC, SAC+HER, and $1:5$ for PPO. We use a batch size of $256$ and a discount factor of $0.99$ for all methods except PPO, for which we use a discount factor of $0.97$. For every environment, we sample evaluation goals from the same distribution as training ones and use a replay buffer of size $10$M for CRL, TD3, TD3+HER, SAC, and SAC+HER. We use $1024$ parallel environments for all methods except for PPO, where we use $4096$ parallel environments to collect data. All experiments are conducted for $50$ million environment steps.
B.3 Benchmark parameters
@@ Appendix F Pseudocode
Pseudocode for the contrastive learning algorithms studied is presented in Algorithm 1.
@@
[…]
3:Observe state $s$ and sample an action $a\sim\pi(s,g)$
4:Execute $a$ in the environment
5:Observe next state $s^{\prime}$ and done signal $d$ to indicate whether $s^{\prime}$ is terminal
[…]
7:Reset environment state and sample new goal
8:Store current trajectory for this environment in $\mathcal{D}$
[…]
10: Randomly sample (with discount) a batch $\mathcal{B}$ from $\mathcal{D}$ of state-action pairs and goals from their future
11: Update critic: $(\phi,\psi)\leftarrow(\phi,\psi)-\alpha\nabla_{\phi,\psi}\bigl{[}\mathcal{L}_{\text{Critic}}(\mathcal{B};\phi,\psi)+\beta\mathcal{L}_{\text{logsumexp}}(\mathcal{B},\phi,\psi)\bigr{]}$
[…]